OpenStudy (anonymous):

What are the exact solutions of x2 + 5x + 1 = 0?

OpenStudy (tyteen4a03):

I've already told you in another question that you need to make use of the Quadratic Formula.

OpenStudy (anonymous):

I don't know how to use it -.-

OpenStudy (tkhunny):

Keren, Keren... You must get it. If you are tasked with finding exact solutions to quadratic equations, you MUST have been exposed to the Quadratic Formula - well, unless you were sleeping in class, or something. It can be developed by applying "Completing the Square" to an arbitrary Quadratic Equation. It is just this: Given a Quadratic Equation: \(a\cdot x^{2} + b\cdot x + c = 0\), both solutions, Real or Complex, are described by \(x = \dfrac{-b\pm\sqrt{b^{2}-4ac}}{2a}\). That's it. In your case, you have a = 1, b = 5, and c = 1. 1) You must see where I got those. 2) You must substitute into the Quadratic Formula and simplify. Now GO! Let's see it.

OpenStudy (anonymous):

It's B